Solve the quadratic equation
5x²-60x+180=0

Answer by nerdybill(5404)

5x²-60x+180=0
Factor out a 5:
5(x²-12x+36)=0
Factoring:
5(x-6)(x-6)=0
.
solution:
x = 6